# A nontrivial uniform algebra regular on the Cantor set

### v1

@@ 1. Introduction
It is well known that every uniform algebra that is regular on its maximal ideal space is normal. The first nontrivial normal uniform algebra was found by Robert McKissick [14]. However, the first example of a nontrivial uniform algebra that is regular on its Shilov boundary appears to be the fiber algebra $A_{\alpha}$ of Kenneth Hoffman and Isadore Singer [10, pp. 187–190] (see also [11, Section 7]). The fiber algebra $A_{\alpha}$ seems also to be the only example in the literature (until now) of a uniform algebra that is regular on its Shilov boundary but is not normal. (Note that a uniform algebra that is regular on its Shilov boundary is normal if and only if its Shilov boundary coincides with its maximal ideal space.) In fact, $A_{\alpha}$ has the stronger property that $|A_{\alpha}|$ ($\{|f|:f\in A_{\alpha}\}$) is normal on the Shilov boundary for $A_{\alpha}$. However, the Shilov boundary for $A_{\alpha}$ is not metrizable. Our examples appear to be the first known examples of uniform algebras $A$ on compact metrizable spaces $X$ such that $A$ is regular on $X$ but $A$ is not normal. Like $A_{\alpha}$, our examples have the property that $|A|$ is normal on $X$.
@@
Since every Dirichlet algebra on a totally disconnected compact space is trivial [10, p. 182], the approach used in this paper will not yield such an algebra. It seems that the only nontrivial Dirichlet algebra known to be regular on a space is a recent example of the second author [13, Theorem 1.7]. That Dirichlet algebra is in fact strongly regular and is defined on a compact metrizable space of topological dimension 1 with trivial Čech cohomology in all degrees. As already mentioned, by a theorem of Wilken there are no nontrivial strongly regular uniform algebras on $[0,1]$.
@@
From Theorem 1.2 we will obtain results for more general compact spaces, provided that they contain a Cantor set.
@@
Let $X$ be a compact space that contains a Cantor set. Then there exists a nontrivial uniform algebra $A$ on $X$ such that $A$ is logmodular and $|A|$ is normal on $X$.
@@
There exists a nontrivial uniform algebra $A$ on $[0,1]$ such that $A$ is logmodular and $|A|$ is normal on $[0,1]$.
@@
Let $X$ be a compact metrizable space without isolated points. Then there exists a nontrivial essential uniform algebra $A$ on $X$ such that $A$ is logmodular and $|A|$ is normal on $X$.
As the above consequences of Theorem 1.2 demonstrate, the Cantor set plays a key role in studying the existence of uniform algebras with particular properties. It is thus of interest to investigate further what properties uniform algebras on the Cantor set can possess. One property of particular interest is the presence of dense squares. The presence of dense squares in a uniform algebra $A$ implies that $A$ has no nontrivial Gleason parts, no nonzero bounded point derivations, and in particular, no analytic discs in its maximal ideal space. As can be seen from the arguments in the last section of our paper, applying a certain form of the Cole construction to any uniform algebra on the Cantor set yields another uniform algebra on the Cantor set that has dense squares. By applying the construction to the uniform algebra of Theorem 1.2 we will obtain the following result.
@@
In Section 2 we recall some standard notation, terminology, and results. In Section 3 we present results that ensure that certain spaces that will arise are free of isolated points. In Section 4 we prove Theorems 1.2 and 1.3. In Section 5 we prove Theorem 1.4, Corollary 1.5, and Theorem 1.6, along with some related results concerning preservation of certain uniform algebra properties by the construction introduced in [7]. In Section 6 we prove that Theorem 1.7 follows from Theorem 1.2 by using Brian Cole’s method of root extensions.
@@ 2. Preliminaries
Throughout the remainder of the section, $A$ will denote a uniform algebra on $X$.
@@
The uniform algebra $A$ is strongly regular at $x\in X$ if each function in $A$ that vanishes at $x$ is a (uniform) limit of functions in $A$ that vanish on a neighborhood of $x$. The uniform algebra $A$ is strongly regular if $A$ is strongly regular at every point of $X$. By a result of Wilken [18, Corollary 1], every strongly regular uniform algebra is normal.
@@
A subset $E$ of $X$ is a set of antisymmetry (for the uniform algebra $A$ on $X$) if every function $f$ in $A$ that is real-valued on $E$ is constant on $E$. If $X$ itself is a set of antisymmetry, then $A$ is said to be an antisymmetric uniform algebra. It is standard that the maximal sets of antisymmetry are closed in $X$, and that they form a partition of $X$. Moreover, if $Y$ is a maximal set of antisymmetry, then $A|Y$ is an antisymmetric uniform algebra on $Y$. By the Bishop-Stone-Weierstrauss theorem (see [2, Theorem 2.7.5] or [16, Theorem 12.1]), if $A$ is nontrivial, then there must be a maximal set of antisymmetry $Y$ such that $A|Y$ is nontrivial.
@@ 3. Ensuring the absence of isolated points
Recall that a space $X$ is said to be scattered if every nonempty subset of $X$ has an isolated point. A famous theorem of Walter Rudin [15] asserts that if $X$ is a compact scattered space, then there are no nontrivial uniform algebras on $X$. Lee Stout [16, p. 119] observed that Rudin’s theorem can be obtained as a consequence of properties of the maximal antisymmetric decomposition. However, Stout used the Shilov idempotent theorem in this deduction, and here we wish to observe that this deep theorem is not needed.
@@ 4. The main construction
The following lemma is standard. See for instance [12, Chapter II, Section 4] or [5, Theorems 1.4.5 and 1.6.5].
@@
We will inductively choose two sequences $({G}_{j})_{j=0}^{\infty}$ and $({H}_{j})_{j=0}^{\infty}$ of countable subsets of $\mathop{\rm inv}\nolimits A$ such that setting,
@@
Let $n$ be a nonnegative integer, and assume for the purpose of induction that we have chosen ${G}_{j}$ and ${H}_{j}$ for all $0\leq j<n$. (Note that when $n=0$, this assumption is vacuous.) Since $A_{n}$ is separable, $X_{A_{n}}$}} is (compact) metrizable, and $^{X}C_{\mathbb{R}}({{X_{A_{n}}}})$ is separable. Since $A$ is logmodular, Lemma 4.1 yields a countable set of functions ${{G}}_{n}$ in $\mathop{\rm inv}\nolimits A$ such that the closure of the set $\{\log|g|:g\in{{G}}_{n}\}$ in $C_{\mathbb{R}}(X)$ contains $^{X}C_{\mathbb{R}}({{X_{A_{n}}}})$. Let J{J} be the countable collection of pairs of disjoint closed subsets of $X_{A_{n}}$}} given by applying Lemma 4.3. Applying Lemma 4.2 then yields the existence of a countable collection K{K} of clopen subsets of $X$ such that for every pair $(K_{1},K_{2})\in{J}$, there exists $K\in{K}$ such that $\pi_{n}^{-1}(K_{1})\subset K$ and $\pi_{n}^{-1}(K_{2})\subset X\setminus K$. Given $K\in{K}$, the function that is 0 on $K$ and 1 on $X\setminus K$ is continuous; since $A$ is logmodular, it follows that there is a function $h_{K}\in\mathop{\rm inv}\nolimits A$ such that $|h_{K}|<3/2$ on $K$ and $|h_{K}|>2$ on $X\setminus K$. We conclude that there is a countable set ${{H}}_{n}$ of functions in $\mathop{\rm inv}\nolimits A$ such that for every function $h\in{H}_{n}$ the image of the function $|h|$ is contained in $0,3/2)\cup(2,\infty)$ and such that for every pair $(C_{1},C_{2})$ of disjoint closed subsets of $X_{A_{n}}$}} there is a function $h\in{H}_{n}$ such that $|h|<3/2$ on $\pi_{n}^{-1}(C_{1})$ and $|h|>2$ on $\pi_{n}^{-1}(C_{2})$. Thus the desired sequences $({G}_{j})_{j=0}^{\infty}$ and $({H}_{j})_{j=0}^{\infty}$ can be inductively chosen.
@@
The proofs of cases (i) and (iii) follow exactly the same pattern. Consider case (i). We again begin by choosing a function $a\in A$ such that $\overline{a}\notin A$, and set $A_{0}=[\{a\}]$. We then inductively choose a single sequence $({F}_{j})_{j=0}^{\infty}$ of countable subsets of $A$ such that setting,
@@
Assume for the purpose of induction that we have chosen ${F}_{j}$ for all $0\leq j<n$. Let J{J} and K{K} be obtained in the same manner as in the proof of case (ii). Since $|A|$ is normal, given $K\in{K}$, there is a function $f_{K}\in A$ such that $f_{K}=0$ on $K$ and $|f_{K}|=1$ on $X\setminus K$. Consequently, there is a countable set ${F}_{n}$ of functions in $A$ such that for every function $f\in{F}_{n}$ the image of the function $|f|$ is contained in $\mathbb{T}\cup\{0\}$ and such that for every pair $(C_{1},C_{2})$ of disjoint closed subsets of $X_{A_{n}}$}} there is a function $f\in{F}_{n}$ such that $f=0$ on $\pi_{n}^{-1}(C_{1})$ and $|f|=1$ on $\pi_{n}^{-1}(C_{2})$. Thus the desired sequence $({F}_{j})_{j=0}^{\infty}$ can be inductively chosen.
@@
As usual we denote by $H^{\infty}(D)$ the algebra of all bounded holomorphic functions on the open unit disc $D$. Let $X$ be the intersection of the fiber of the maximal ideal space of $H^{\infty}(D)$ over the point $1$ with the Shilov boundary $S$ for $H^{\infty}(D)$. It is well-known that $S$, and hence $X$, is totally disconnected [10, pp. 169-175]. Let $A$ be the algebra obtained by restricting the Gelfand transforms of the functions in $H^{\infty}(D)$ to $X$. Then as shown in [10, pp. 189–190] (see also [11]), $A$ is a uniform algebra on $X$ that has no nontrivial idempotents, $X$ is the Shilov boundary for $A$, and $|A|$ is normal on $X$. (The algebra $A$ is the Hoffman-Singer fiber algebra $A_{\alpha}$ with $\alpha=1$.) It is well known that $H^{\infty}(D)$ is a logmodular algebra [10, p. 182], and it follows that $A$ is logmodular on $X$. Let $\widetilde{B}$ be the uniform algebra on $X_{B}$ obtained by applying Theorem 1.3 to $A$. Then $X_{B}$ is compact, totally disconnected, and metrizable. Since $X_{B}$ is the Shilov boundary for $\widetilde{B}$, as noted above, by Corollary 3.3, $X_{B}$ has no isolated points. Thus $X_{B}$ is a Cantor set, and $\widetilde{B}$ is a uniform algebra with all the required properties. ∎
It is standard that there is a function in $H^{\infty}(D)$ whose real part extends continuously to the boundary of $D$ but whose imaginary part does not [8, p. 377]. It follows that the Hoffman-Singer fiber algebra $A_{1}$ is not antisymmetric. Consequently, the algebra constructed in the proof of Theorem 1.2 can also fail to be antisymmetric (if, for instance, one chooses the sequence $({G}_{j})_{j=0}^{\infty}$ in the proof of Theorem 1.3 to include a nonconstant real-valued function). However, by applying Corollary 3.5, we can obtain an antisymmetric uniform algebra with the properties in Theorem 1.2. Note also that invoking Corollary 3.5 eliminates the need to prove that the space $X_{B}$ in the proof above has no isolated points.
@@ 5. Uniform algebras logmodular and regular on general spaces
We first show that $|A|$ is normal. Consider disjoint closed subsets $K_{0}$ and $K_{1}$ in $X$. By the normality of $|B|$, there is a function $b\in B$ such that $b=0$ on $K_{0}\cap K$ and $|b|=1$ on $K_{1}\cap K$. As a closed subset of a Cantor set, $K_{1}\cap K$ is simply coconnected. Therefore, there is a continuous real-valued function $\sigma:K_{1}\cap K\rightarrow\mathbb{R}$ such that $b=e^{i\sigma}$ on $K_{1}\cap K$. By the Tietze extension theorem, $\sigma$ extends to a continuous function $\widetilde{\sigma}:K_{1}\rightarrow\mathbb{R}$. There is a well-defined continuous function $a:K_{0}\cup K_{1}\cup K\rightarrow\mathbb{C}$ given
@@
First suppose that $x$ lies in none of the sets $K_{n}$ that intersect $E$. (The point $x$ may or may not lie in some $K_{n}$ that is disjoint from $E$.) In this case, let $Z$ denote the quotient space obtained by identifying each $K_{n}$ for $n\in\mathbb{N}$ to a point, and let $p:X\rightarrow Z$ be the canonical quotient map. The space $Z$ is compact and metrizable [7, Lemma 2.6]. Then $p(x)\notin p(E)$, so Urysohn’s lemma provides a function $f^{*}\in C_{\mathbb{R}}(Z)$ such that $f^{*}=1$ on $p(E)$ and $f^{*}=0$ on a neighborhood of $p(x)$. The function $f=f^{*}\circ p$ is in $C_{\mathbb{R}}(X)$ and $f$ is constant on each $K_{n}$, so $f\in A_{X}$. Clearly $f$ has the required properties.
@@
The proof of this theorem will use the following elementary lemma whose proof is given in [13, Lemma 6.3].
@@
The proof is similar to the proof that $A_{X}$ is Dirichlet on $X$ if and only if $A$ is Dirichlet on $K$ given in [13, Theorem 6.2].
@@
Let $Z$ denote the quotient space obtained by identifying each $K_{n}$ for $n>N$ to a point, and let $p:X\rightarrow Z$ be the canonical quotient map. The compact space $Z$ is metrizable [7, Lemma 2.6].

### v2

@@ 1. Introduction
It is well known that every uniform algebra that is regular on its maximal ideal space is normal. The first nontrivial normal uniform algebra was found by Robert McKissick [15]. However, the first example of a nontrivial uniform algebra that is regular on its Shilov boundary appears to be the fiber algebra $A_{\alpha}$ of Kenneth Hoffman and Isadore Singer [11, pp. 187–190] (see also [12, Section 7]). The fiber algebra $A_{\alpha}$ seems also to be the only example in the literature (until now) of a uniform algebra that is regular on its Shilov boundary but is not normal. (Note that a uniform algebra that is regular on its Shilov boundary is normal if and only if its Shilov boundary coincides with its maximal ideal space.) In fact, $A_{\alpha}$ has the stronger property that $|A_{\alpha}|$ ($\{|f|:f\in A_{\alpha}\}$) is normal on the Shilov boundary for $A_{\alpha}$. However, the Shilov boundary for $A_{\alpha}$ is not metrizable. Our examples appear to be the first known examples of uniform algebras $A$ on compact metrizable spaces $X$ such that $A$ is regular on $X$ but $A$ is not normal. Like $A_{\alpha}$, our examples have the property that $|A|$ is normal on $X$.
@@
Since every Dirichlet algebra on a totally disconnected compact space is trivial [11, p. 182], the approach used in this paper will not yield such an algebra. It seems that the only construction of a nontrivial Dirichlet algebra known to be regular on a space is the one recently given by the second author in [14, Theorem 1.7]. The Dirichlet algebra given there is in fact strongly regular and is defined on a compact metrizable space of topological dimension 1 with trivial Čech cohomology in all degrees. As already mentioned, by a theorem of Wilken there are no nontrivial strongly regular uniform algebras on $[0,1]$.
@@
In connection with this theorem, note that no nontrivial uniform algebra on a totally disconnected space can be normal. This follows immediately from the definition of normality for a uniform algebra and the Stone-Weierstrass theorem since for every pair of distinct points $p$ and $q$ in a totally disconnected compact space there is a separation $P,Q$ of the space with $p\in P$ and $q\in Q$.
@@
From Theorem 1.2 we will obtain results for more general compact spaces, provided that they contain a Cantor set. The uniform algebras in each of these theorems are regular on the given space but are not normal. We remind the reader that there are no nontrivial normal uniform algebras on a totally disconnected space. As noted earlier, the existence of nontrivial normal Dirichlet algebras is proven in [14].
@@
Let $X$ be a compact space that contains a Cantor set. Then there exists a nontrivial uniform algebra $A$ on $X$ such that $A$ is logmodular and $|A|$ is normal on $X$ but $A$ is not normal.
@@
There exists a nontrivial uniform algebra $A$ on $[0,1]$ such that $A$ is logmodular and $|A|$ is normal on $[0,1]$ but $A$ is not normal.
@@
Let $X$ be a compact metrizable space without isolated points. Then there exists a nontrivial essential uniform algebra $A$ on $X$ such that $A$ is logmodular and $|A|$ is normal on $X$ but $A$ is not normal.
As the above consequences of Theorem 1.2 demonstrate, the Cantor set plays a key role in studying the existence of uniform algebras with particular properties. It is thus of interest to investigate further what properties uniform algebras on the Cantor set can possess. One property of particular interest is the presence of dense squares. The presence of dense squares in a uniform algebra $A$ implies that $A$ has no nontrivial Gleason parts, no nonzero bounded point derivations, and in particular, no analytic discs in its maximal ideal space. A theorem of E. M. Chirka asserts that there are no nontrivial uniform algebras with dense squares on locally connected spaces [17, Theorem 13.15 and Lemma 13.16] (see also [4]). In contrast, a construction of Brian Cole yields uniform algebras that do have dense squares (or even every element a square) on certain nonlocally connected spaces. In general though, as remarked by Lee Stout [17, p. 200] “it is not at all evident what topological properties” these spaces have. However, it can be seen from the arguments in the last section of our paper that applying a certain form of the Cole construction to any uniform algebra on the Cantor set yields another uniform algebra on the Cantor set that has dense squares. By applying the construction to the uniform algebra of Theorem 1.2 we will obtain the following result.
@@
In Section 2 we recall some standard notation, terminology, and results. In Section 3 we present results that ensure that certain spaces that will arise are free of isolated points. In Section 4 we prove Theorems 1.2 and 1.3. In Section 5 we prove Theorem 1.4, Corollary 1.5, and Theorem 1.6, along with some related results concerning preservation of certain uniform algebra properties by the construction introduced in [8]. In Section 6 we prove that Theorem 1.7 follows from Theorem 1.2 by using Cole’s method of root extensions.
@@ 2. Preliminaries
Throughout the remainder of this section, $A$ will denote a uniform algebra on $X$.
@@
The uniform algebra $A$ is strongly regular at $x\in X$ if each function in $A$ that vanishes at $x$ is a (uniform) limit of functions in $A$ that vanish on a neighborhood of $x$. The uniform algebra $A$ is strongly regular if $A$ is strongly regular at every point of $X$. By a result of Wilken [19, Corollary 1], every strongly regular uniform algebra is normal.
@@
A subset $E$ of $X$ is a set of antisymmetry (for the uniform algebra $A$ on $X$) if every function $f$ in $A$ that is real-valued on $E$ is constant on $E$. If $X$ itself is a set of antisymmetry, then $A$ is said to be an antisymmetric uniform algebra. It is standard that the maximal sets of antisymmetry are closed in $X$, and that they form a partition of $X$. Moreover, if $Y$ is a maximal set of antisymmetry, then $A|Y$ is an antisymmetric uniform algebra on $Y$. By the Bishop-Stone-Weierstrauss theorem (see [2, Theorem 2.7.5] or [17, Theorem 12.1]), if $A$ is nontrivial, then there must be a maximal set of antisymmetry $Y$ such that $A|Y$ is nontrivial.
@@ 3. Ensuring the absence of isolated points
Recall that a space $X$ is said to be scattered if every nonempty subset of $X$ has an isolated point. A famous theorem of Walter Rudin [16] asserts that if $X$ is a compact scattered space, then there are no nontrivial uniform algebras on $X$. Stout [17, p. 119] observed that Rudin’s theorem can be obtained as a consequence of properties of the maximal antisymmetric decomposition. However, Stout used the Shilov idempotent theorem in this deduction, and here we wish to observe that this deep theorem is not needed.
@@ 4. The main construction
The following lemma is standard. See for instance [13, Chapter II, Section 4] or [6, Theorems 1.4.5 and 1.6.5].
@@
We will inductively choose two sequences $({G}_{j})_{j=0}^{\infty}$ and $({H}_{j})_{j=0}^{\infty}$ of countable subsets of $\mathop{\rm inv}\nolimits A$ such that setting
@@
Let $n$ be a nonnegative integer, and assume for the purpose of induction that we have chosen ${G}_{j}$ and ${H}_{j}$ for all $0\leq j<n$. (When $n=0$, this assumption is vacuous.) Note that $A_{n}$ is separable since it is generated by a countable set of functions. Therefore, $X_{A_{n}}$}} is (compact) metrizable, and $^{X}C_{\mathbb{R}}({{X_{A_{n}}}})$ is separable. Since $A$ is logmodular, Lemma 4.1 yields a countable set of functions ${{G}}_{n}$ in $\mathop{\rm inv}\nolimits A$ such that the closure of the set $\{\log|g|:g\in{{G}}_{n}\}$ in $C_{\mathbb{R}}(X)$ contains $^{X}C_{\mathbb{R}}({{X_{A_{n}}}})$. Let J{J} be the countable collection of pairs of disjoint closed subsets of $X_{A_{n}}$}} given by applying Lemma 4.3. Applying Lemma 4.2 then yields the existence of a countable collection K{K} of clopen subsets of $X$ such that for every pair $(K_{1},K_{2})\in{J}$, there exists $K\in{K}$ such that $\pi_{n}^{-1}(K_{1})\subset K$ and $\pi_{n}^{-1}(K_{2})\subset X\setminus K$. Given $K\in{K}$, the function that is 0 on $K$ and 1 on $X\setminus K$ is continuous; since $A$ is logmodular, it follows that there is a function $h_{K}\in\mathop{\rm inv}\nolimits A$ such that $|h_{K}|<3/2$ on $K$ and $|h_{K}|>2$ on $X\setminus K$. We conclude that there is a countable set ${{H}}_{n}$ of functions in $\mathop{\rm inv}\nolimits A$ such that for every function $h\in{H}_{n}$ the image of the function $|h|$ is contained in $0,3/2)\cup(2,\infty)$ and such that for every pair $(C_{1},C_{2})$ of disjoint closed subsets of $X_{A_{n}}$}} there is a function $h\in{H}_{n}$ such that $|h|<3/2$ on $\pi_{n}^{-1}(C_{1})$ and $|h|>2$ on $\pi_{n}^{-1}(C_{2})$. Thus the desired sequences $({G}_{j})_{j=0}^{\infty}$ and $({H}_{j})_{j=0}^{\infty}$ can be inductively chosen.
@@
The proofs of cases (i) and (iii) follow exactly the same pattern. Consider case (i). We again begin by choosing a function $a\in A$ such that $\overline{a}\notin A$, and set $A_{0}=[\{a\}]$. We then inductively choose a single sequence $({F}_{j})_{j=0}^{\infty}$ of countable subsets of $A$ such that setting
@@
Assume for the purpose of induction that we have chosen ${F}_{j}$ for all $0\leq j<n$. Let J{J} and K{K} be obtained in the same manner as in the proof of case (ii). Since $|A|$ is normal on $X$, given $K\in{K}$, there is a function $f_{K}\in A$ such that $f_{K}=0$ on $K$ and $|f_{K}|=1$ on $X\setminus K$. Consequently, there is a countable set ${F}_{n}$ of functions in $A$ such that for every function $f\in{F}_{n}$ the image of the function $|f|$ is contained in $\{0,1\}$ and such that for every pair $(C_{1},C_{2})$ of disjoint closed subsets of $X_{A_{n}}$}} there is a function $f\in{F}_{n}$ such that $f=0$ on $\pi_{n}^{-1}(C_{1})$ and $|f|=1$ on $\pi_{n}^{-1}(C_{2})$. Thus the desired sequence $({F}_{j})_{j=0}^{\infty}$ can be inductively chosen.
@@
As usual we denote by $H^{\infty}(D)$ the algebra of all bounded holomorphic functions on the open unit disc $D$. Let $X$ be the intersection of the fiber of the maximal ideal space of $H^{\infty}(D)$ over the point $1$ with the Shilov boundary $S$ for $H^{\infty}(D)$. It is well known that $S$, and hence $X$, is totally disconnected [11, pp. 169-175]. Let $A$ be the algebra obtained by restricting the Gelfand transforms of the functions in $H^{\infty}(D)$ to $X$. Then as shown in [11, pp. 189–190] (see also [12]), $A$ is a uniform algebra on $X$ that has no nontrivial idempotents, $X$ is the Shilov boundary for $A$, and $|A|$ is normal on $X$. (The algebra $A$ is the Hoffman-Singer fiber algebra $A_{\alpha}$ with $\alpha=1$.) It is well known that $H^{\infty}(D)$ is a logmodular algebra [11, p. 182], and it follows that $A$ is logmodular on $X$. Let $\widetilde{B}$ be the uniform algebra on $X_{B}$ obtained by applying Theorem 1.3 to $A$. Then $X_{B}$ is compact, totally disconnected, and metrizable. Since $X_{B}$ is the Shilov boundary for $\widetilde{B}$, as noted above, by Corollary 3.3, $X_{B}$ has no isolated points. Thus $X_{B}$ is a Cantor set, and $\widetilde{B}$ is a uniform algebra with all the required properties. ∎
It is standard that there is a function in $H^{\infty}(D)$ whose real part extends continuously to the boundary of $D$ but whose imaginary part does not [9, p. 377]. It follows that the Hoffman-Singer fiber algebra $A_{1}$ is not antisymmetric. Consequently, the algebra constructed in the proof of Theorem 1.2 can also fail to be antisymmetric (if, for instance, one chooses the sequence $({G}_{j})_{j=0}^{\infty}$ in the proof of Theorem 1.3 to include a nonconstant real-valued function). However, by applying Corollary 3.5, we can obtain an antisymmetric uniform algebra with the properties in Theorem 1.2. Note also that invoking Corollary 3.5 eliminates the need to prove that the space $X_{B}$ in the proof above has no isolated points.
@@ 5. Uniform algebras logmodular and regular on general spaces
We first show that $|A|$ is normal on $X$. Consider disjoint closed subsets $K_{0}$ and $K_{1}$ in $X$. By the normality of $|B|$, there is a function $b\in B$ such that $b=0$ on $K_{0}\cap K$ and $|b|=1$ on $K_{1}\cap K$. As a closed subset of a Cantor set, $K_{1}\cap K$ is simply coconnected. Therefore, there is a continuous real-valued function $\sigma:K_{1}\cap K\rightarrow\mathbb{R}$ such that $b=e^{i\sigma}$ on $K_{1}\cap K$. By the Tietze extension theorem, $\sigma$ extends to a continuous function $\widetilde{\sigma}:K_{1}\rightarrow\mathbb{R}$. There is a well-defined continuous function $a:K_{0}\cup K_{1}\cup K\rightarrow\mathbb{C}$ given
@@
The uniform algebra $A$ is not normal since its restriction to the Cantor set $K$ is the nontrivial uniform algebra $B$. ∎
@@
First suppose that $x$ lies in none of the sets $K_{n}$ that intersect $E$. (The point $x$ may or may not lie in some $K_{n}$ that is disjoint from $E$.) In this case, let $Z$ denote the quotient space obtained by identifying each $K_{n}$ for $n\in\mathbb{N}$ to a point, and let $p:X\rightarrow Z$ be the canonical quotient map. The space $Z$ is compact and metrizable [8, Lemma 2.6]. Then $p(x)\notin p(E)$, so Urysohn’s lemma provides a function $f^{*}\in C_{\mathbb{R}}(Z)$ such that $f^{*}=1$ on $p(E)$ and $f^{*}=0$ on a neighborhood of $p(x)$. The function $f=f^{*}\circ p$ is in $C_{\mathbb{R}}(X)$ and $f$ is constant on each $K_{n}$, so $f\in A_{X}$. Clearly $f$ has the required properties.
@@
The proof of this theorem will use the following elementary lemma whose proof is given in [14, Lemma 6.3].
@@
The proof is similar to the proof that $A_{X}$ is Dirichlet on $X$ if and only if $A$ is Dirichlet on $K$ given in [14, Theorem 6.2].
@@
Let $Z$ denote the quotient space obtained by identifying each $K_{n}$ for $n>N$ to a point, and let $p:X\rightarrow Z$ be the canonical quotient map. The compact space $Z$ is metrizable [8, Lemma 2.6].